OpenStudy (anonymous):

what are the equivalent measurements???? 4/8 inch = ?/16 inch 1 inch = ?/16 inch

OpenStudy (anonymous):

1 1/2 inch = 1 ?/16

OpenStudy (anonymous):

4/8=8/16 1=16/16 1 1/2=1 8/16

OpenStudy (anonymous):

x = 16 *4/8 = 8 in x = 16 * 1 = 16 in